OpenStudy (anonymous):

Simplify -1/x+2 + x-1/4/ x+3/x+2

OpenStudy (anonymous):

is it x-1/4 + x+3/x+2?

OpenStudy (anonymous):

or are there fractions in the fraction?

OpenStudy (anonymous):

\[\frac{ \frac{ \frac{ -1 }{ x+2 }+\frac{ x-1 }{ 4 } }{ x+3 } }{ x+2 }\]

OpenStudy (anonymous):

\[=\frac{ \frac{ -4+(x-1)(x+2) }{ 4(x+2) } }{ \frac{ x+3 }{ x+2 } }\] \[=\frac{ -4+x^2+2x-x-2 }{ 4(x+2) }\times \frac{ x+2 }{ x+3 }\] \[\frac{ x^2+x-6 }{ 4 }\times \frac{ 1 }{ x+3 }\] \[\frac{ (x+3)(x-2) }{ 4 }\times \frac{ 1 }{ x+3 }\]

OpenStudy (anonymous):

\[\frac{ x-2 }{ 4 }\]

OpenStudy (anonymous):

Missed a few equal signs but you can add them yourselves.

OpenStudy (anonymous):

Thank you so much!